# An International Agreement to Prevent the Premature Creation of Artificial Superintelligence

### v1

@@ Abstract
Many experts argue that artificial superintelligence (ASI), if developed prematurely, poses catastrophic risks including human extinction from misaligned ASI, geopolitical instability, and misuse by malicious actors. This report proposes an international agreement to prevent the premature development of ASI until AI development can proceed without these risks. The agreement halts dangerous AI capabilities advancement while preserving access to current, beneficial AI applications.
The proposed framework centers on a coalition led by the United States and China that would restrict the scale of AI training and dangerous AI research. Due to the lack of trust between relevant parties, verification is a key part of the agreement. Limits on the scale of AI training are operationalized by FLOP thresholds and verified through the tracking of AI chips and verification of chip use. Dangerous AI research—that which advances toward artificial superintelligence or endangers the agreement’s verifiability—is stopped via legal prohibitions and multifaceted verification.
We believe the proposal would be technically sufficient if implemented today, but advancements in AI capabilities or development methods would hurt its efficacy. Simultaneously, there is not yet political will for such an agreement. Despite these concerns, we hope this agreement can provide direction for AI governance research and policy.
@@ 1 Introduction
The 2025 book If Anyone Builds It, Everyone Dies argues that the development of artificial superintelligence (ASI, AI smarter than humans at all cognitive tasks) “using anything remotely like current techniques, based on anything remotely like the present understanding of AI” leads to human extinction (yudkowsky2025ifanyone). If that were true, what should we do? Even if there were “only” a 10% chance of human extinction, as some other leading experts claim (pdoom2025wikipedia), how should governments respond?
In this report, we propose an international agreement which, if implemented, would postpone the development of ASI until proper mitigation efforts are developed and implemented (i.e., potentially decades). Appendix A lays out draft text for such an agreement. Practically speaking, this agreement is not going to be implemented tomorrow, but we hope it can serve as an end goal for AI governance, and in Appendix C we lay out a staged implementation that iteratively builds up to the final agreement.
@@
  • We briefly overview AI risks and the key details of the strategic situation informing our proposal.
@@
  • We provide the full text of an example agreement, detailed commentary, and relation to existing agreements in Appendix A. We discuss immediate steps different actors could take toward this agreement in Appendix B. We explain a staged implementation of the agreement that slowly builds from the current world to the agreement in Appendix C. We detail various approaches that could help locate existing AI chips and bring them under monitoring in Appendix D.
@@ 2 An overview of AI risks
There are many large-scale risks from advanced AI (i.e. AI much more powerful than today’s AI technology) that might warrant slowing or stopping AI progress for a significant period (e.g., decades). Most of these risks stem from predictions that AI becomes an increasingly capable tool: Rapid automation may destabilize economies (drago2025intelligence; erdil2025gate). Capabilities that undermine strategic deterrence and incentivize first strikes may trigger wars and conflicts (PE-A3691-4; hendrycks_superintelligence_2025). Malicious actors may misuse AI to engage in widescale cyberattacks or develop novel biological weapons (zelikow2024defense; crawford2024securing; esvelt2022delay). AI used for mass surveillance, propaganda, and police-state oppression may enable extreme concentrations of power (davidson2025aienabledcoupsho; bullock2025agi; drago2025intelligence).
@@
But while the above risks may lead to unprecedented catastrophes, humanity can probably survive them. A greater and more irreversible risk stems from AIs that don’t remain passive tools, while becoming capable enough to permanently disempower humanity. This plausibly leads to human extinction, either as an intentional act by the AIs (e.g., to prevent humans from making powerful competing AIs) or as a side effect of optimizing for goals that don’t specifically designate human survival (yudkowsky2022agi; ngo_alignment_2025).
@@
This paper focuses on an international agreement to forestall ASI development until it can be done safely. A halt like this is primarily motivated by concerns about misalignment and AI takeover, and we think it is the only acceptable mitigation to those risks. But a halt would also be an effective tool for many other risks, such as geopolitical destabilization, mass AI-caused unemployment, and misuse of powerful AIs (barnett2025ai). Therefore, others may support a halt even if they are not concerned with misalignment.
@@
The deep learning paradigm underpinning modern AI development seems highly prone to producing agents that are not aligned with humanity’s interests. Under this paradigm, AI behavior is not hand-coded by human engineers, but instead emerges from automated training processes. Rather than producing lines of code that can be debugged when something goes wrong, these processes configure hundreds of billions to trillions of model weights that direct the flow of information through the AI. While it is sometimes possible to identify and activate features of these weights correlated with certain inputs or outputs, the deep machinery of these minds is effectively opaque to human observers. No one can say with any certainty why an AI made a particular choice or how it would act in situations that differ from its training.
@@
Consequently, no one can manually adjust an AI’s weights to give it a particular goal or set of preferences. And yet, as seen in the growing ability of AIs to complete long tasks, AIs are becoming increasingly capable of pursuing goals, for instance, in the setting of agentic software engineering (kwa2025measuring). Various well-documented examples of misbehavior by these same models (both in the lab and in the wild (baker2025monitoring; bondarenko2025demonstrating; lynch2025agentic; openai2025sycophancy; metr2025reward)) demonstrate that today’s training methods do not produce robustly aligned agents that could be trusted not to disempower humanity if they had the capacity.
@@
It is hard to predict when AI labs might succeed at what is frequently their stated goal: artificial superintelligence (ASI) that can surpass humans at all cognitive tasks. That said, many predict ASI could be developed soon. The field has seen enormous progress in just three years—from AIs that couldn’t hold a conversation to “reasoning models” that solve hard exam questions (EpochLLMBenchmarkingHub2024) and that some major tech companies claim write a substantial fraction of their code (chandra2024ai; novet2025satya). Another few years of such progress might well push capabilities well beyond human level in all or nearly all domains.
@@
One accelerating factor complicating all forecasts is the possibility of AIs automating the process of AI research and development itself (davidson2025threetypesofinte; eth2025willairdautomati; wijk_re-bench_2024). How close to this threshold they might be is unclear, but some independent experts have given median prediction dates ranging from 2028–2033 (lifland2025timelines). If or when this threshold is crossed, a feedback loop where AI R&D is conducted by AI might rapidly—in months or a couple years—produce much more capable AIs, including ASI.
@@
On the current trajectory, the core challenge of AI alignment is unlikely to be solved before the first ASIs are built. Evidence for this can be seen in the failure to solve much easier related problems. Today’s frontier models are not robust to jailbreaks: It is relatively easy to craft inputs that lead to outputs that AI developers attempt to avoid (andriushchenko2024agentharm; andriushchenko2024jailbreaking; zou2023universal; anil_many-shot_2024; wang_jailbreak_2024). Current models, especially reasoning models, display a strong propensity to reward hack (exploiting unintended solutions) and lie to their users, against the best efforts of their developers (bondarenko2025demonstrating; wijk_re-bench_2024; chowdhury2025investigating; anthropic2025claude37; baker2025monitoring). Far tougher alignment challenges are likely to appear only at advanced capability levels where humanity would be unable to recover from failures (yudkowsky2022agi).
@@
There is a broad consensus in the field that catastrophe resulting from misaligned AI is a significant threat that should be addressed (bengio2024managing). Hundreds of luminaries signed on to a 2023 statement that “Mitigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war.” (cais_statement_2023) That same year, Yoshua Bengio, a Turing Award recipient and the most cited living scientist, said he thinks there is a (lavoipierre2023ai), “twenty percent probability that it turns out catastrophic.” In a 2023 survey (grace2025thousands), 38% of survey respondents at top AI conferences said there was at least a 10% chance that AI’s outcome will be at least as bad as human extinction.
@@
This understanding extends to AI company leaders themselves: Anthropic CEO Dario Amodei has predicted a 10-25% (bartlett2023anthropic) that “something goes really quite catastrophically wrong on the scale of human civilization.” OpenAI’s Sam Altman has called superintelligence “probably the greatest threat to the continued existence of humanity” (altman2015machine), and xAI’s Elon Musk has said, “AI is a fundamental risk to the existence of human civilization”  (london2017musk). Musk has indicated that his participation in this ASI race stems from the belief that it would happen even without him (musk2025interview). Those racing toward superintelligence, to the extent they are concerned with catastrophic risks, are stuck in a collective action problem (armstrong2016racing). Halting the race and avoiding extinction would require global coordination encompassing all relevant companies and governments.
@@ 3 The strategic situation
A 10% chance of extinction is unacceptable (and some experts put the likelihood much higher (pdoom2025wikipedia)). To mitigate these risks, various approaches have been proposed: Have governments tightly regulate and license AI development, empower defensive actors early (buterin_my_2023; nielsen_notes_2024; bernardi_societal_2025), implement a joint global AI development project to avoid racing (brundage_my_2024; hausenloy_multinational_2023), race ahead while slowing opponents (aschenbrenner_situational_2024; amodei_machines_2024), or prevent the creation of this technology before we are ready (barnett2025ai; ramiah2025toward; aguirre2025keep), to name a few.
@@
We think the only acceptable solution is to prevent the creation of ASI until catastrophic risks are mitigated. The state of AI alignment research is too nascent, and the problem too difficult. Nobody knows how to align advanced AIs to human values, including regulators or an international standards body (and we address some of these alternative plans in more depth in Section 5). Therefore, we think the right approach to this problem is to forestall the development of ASI for as long as is necessary to solve critical alignment challenges. An agreement establishing such a halt would need to be globally enforced, verifiable, and ready to implement once political will is achieved.
@@
For a halt to be effective, it must extend to all countries. If only cautious actors stop, other parties will continue the march toward superintelligence. The United States (U.S.) pausing on its own might buy around six months until the nearest company in the People’s Republic of China (PRC) catches up, and a U.S.–PRC agreement without global enforcement would similarly fall to the next closest actor, likely in a couple years (artificialanalysis2025countries) (though it is unclear how long it would take to get from the current frontier to ASI). The U.S. and PRC are the preeminent superpowers and the centers of most frontier AI development; therefore, they are essential stakeholders in such a deal whose strategic interests must be accommodated. Due to their influence, they are likely leaders in future AI agreements, even if other countries participate as well.
@@
A coalition starting with the U.S. and PRC would likely be sufficient to govern AI development around the world, given these countries’ geopolitical influence, in addition to their current positions as AI leaders. We expect other countries to sign on to such an agreement once they understand AI risks better and conclude that international coordination is necessary to avoid these risks. Participation could also be encouraged with benefits such as security and access to AI infrastructure for allowed use. Where stronger incentives prove necessary, standard tools such as economic sanctions and visa bans may apply. We believe a very ASI-concerned coalition headed by the U.S. and PRC would be sufficient to integrate other potential AI developers into the agreement or to sufficiently suppress other AI development so that it does not pose a threat.
@@
This paper describes an agreement headed by the U.S. and PRC and potentially other key states as appropriate to the current geopolitical reality. All parties will desire assurance that other parties are not engaged in secret AI capabilities projects. Given a low-trust environment, parties must have the means to verify compliance to their satisfaction, and our agreement facilitates such verification. It gives special consideration to the independent verification efforts of the U.S. and the PRC.
@@
There remain significant political obstacles to such an agreement occurring today. Nevertheless, to accept even a 10% chance of extinction is wholly inconsistent with how we manage other risks; we routinely adopt stringent mitigations for risks of far lower likelihood and consequence in national security, such as in the nuclear industry, and in many other fields (koessler2024risk; nrc2017approach).
@@ 4 The Agreement
AI development in the current deep learning paradigm is driven both by the large-scale employment of specialized computer chips and by improvements in AI algorithms (epoch2024trainingcomputeoffrontieraimodelsgrowsby45xperyear; kaplan_scaling_2020; ho_algorithmic_2024). Therefore, the coalition stops the advancement of AI capabilities toward superintelligence by restricting the scale of AI training (operationalized as total training FLOP) and by restricting research into AI algorithms and other areas that would undermine the governance regime.
@@
AI training runs above the Strict Threshold (e.g., $10^{24}$ FLOP) are prohibited. Training runs below this threshold but above the Monitored Threshold (e.g., $10^{22}$ FLOP) must be approved and monitored by coalition authorities. Training runs below the Monitored Threshold can take place without approval or monitoring. Verification of these restrictions is practical because AI chips are expensive and specialized, and thousands of them are needed for frontier AI development. This approach helps ensure that training runs that advance toward ASI are prevented while still allowing continued use of current AI technology.
@@
Figure 2: The training compute used to train various notable AI models in the last few years, along with the two thresholds in this agreement. In the agreement, new training above $10^{24}$ FLOP would be prohibited and new training between $10^{22}$ and $10^{24}$ FLOP would be monitored. Due to a lack of public data, only some models have confident estimates for the training compute used (we only label models with confident compute estimates). The number in parentheses is the model’s score on the Artificial Analysis Intelligence Index. Data is based on EpochAIModels2025 and artificialanalysis_models. Cost estimates are based on October 2025 rental prices for B200 GPUs from CoreWeave, assuming 50% utilization in FP8.
@@
To achieve this verification, the coalition locates existing AI chips, tracks new AI chips being produced, and monitors AI chip use in order to ensure chips are not being used for restricted activities (scher_mechanisms_2024; baker2025verifying; harack2025verification). Chips are located through coordinated efforts by intelligence agencies and domestic law-enforcement agencies from coalition countries. These efforts involve combining supply chain tracking, mandatory reporting, state intelligence gathering, power consumption monitoring, challenge inspections, whistleblowers, and more. Appendix D provides more detail on how this could be done. Tracking new AI chips relies on the concentrated chip supply chain and serves the goal of making sure chips are not smuggled outside of the monitoring regime. Finally, the chips themselves are monitored in order to verify that they are only being used for acceptable use cases.
@@
AI chip monitoring starts with discerning whether the chips are being used for inference on existing AIs or training of new AIs. The coalition works to develop tamper-resistant on-chip mechanisms for such purposes, as finer-grained control permits greater use of AI chips for safe applications, with fewer intrusive inspections. But initially inspectors are given ongoing physical access to chips, as is likely needed for robust chip-use verification (aarne2024secure; kulp2024hardwareenabled). To ensure chip use verification is applied, the coalition prohibits large concentrations of chips (e.g., greater than 16 H100-equivalents; 16 H100s cost approximately \$500,000 USD in 2025) outside of monitored facilities.
@@
Despite prohibitions on large-scale training, AIs could continue getting more capable via improvements to the algorithms and software used to train them. Therefore, the coalition adopts appropriate restrictions on research that contributes to frontier AI development or research that could endanger the verification methods in the agreement. Coalition members draw on their experience restricting research in other dangerous fields such as nuclear and chemical weapons, working to keep restrictions as narrow as possible. The coalition makes explicit carve-outs for safe, narrow AI activities, such as self-driving cars and other uses that provide benefits to society.
@@
Verification that members are adhering to research restrictions is aided by the fact that relatively few people have the skills to contribute to this research—likely only thousands or tens of thousands, see Appendix A, Article IX. Nations verify compliance by way of intelligence gathering, interviews with the researchers, whistleblower programs, and more. Additionally, inspectors verify that the medium-scale training allowed by the agreement uses only approved methods (e.g., from a Whitelist) and does not make use of any novel AI methods or algorithms; these would be evidence that restricted research had taken place.
@@
Should nonproliferation fail, or a rogue actor is found to be working towards ASI, coalition members can employ their agreed upon enforcement mechanisms. These start with standard tools of international pressure such as diplomacy, economic sanctions, and visa bans. Depending on the nature of the developments, countries could escalate to other means at their disposal to disrupt, slow, or impede ASI development. This draws parallels to other contexts where countries have intervened to stop violations of arms-control and nonproliferation agreements.

### v2

@@ Abstract
Many experts argue that premature development of artificial superintelligence (ASI) poses catastrophic risks, including the risk of human extinction from misaligned ASI, geopolitical instability, and misuse by malicious actors. This report proposes an international agreement to prevent the premature development of ASI until AI development can proceed without these risks. The agreement halts dangerous AI capabilities advancement while preserving access to current, safe AI applications.
The proposed framework centers on a coalition led by the United States and China that would restrict the scale of AI training and dangerous AI research. Due to the lack of trust between parties, verification is a key part of the agreement. Limits on the scale of AI training are operationalized by FLOP thresholds and verified through the tracking of AI chips and verification of chip use. Dangerous AI research—that which advances toward artificial superintelligence or endangers the agreement’s verifiability—is stopped via legal prohibitions and multifaceted verification.
We believe the proposal would be technically sufficient to forestall the development of ASI if implemented today, but advancements in AI capabilities or development methods could hurt its efficacy. Additionally, there does not yet exist the political will to put such an agreement in place. Despite these challenges, we hope this agreement can provide direction for AI governance research and policy.
@@ 1 Introduction
The 2025 book If Anyone Builds It, Everyone Dies argues that the development of artificial superintelligence (ASI, AI substantially smarter than humans at all cognitive tasks) “using anything remotely like current techniques, based on anything remotely like the present understanding of AI” would lead to human extinction (yudkowsky2025ifanyone). If that’s the case, what should we do about it? Experts differ in their assessment of the situation. Many think that there is around a 10% chance of future AI advances causing human extinction (for aggregations of expert opinion see grace2025thousands; pdoom2025wikipedia; controlai_quotes_2025). Even if there were “only” a 10% chance of human extinction, how should governments respond?
In this report, we propose an international agreement which, if implemented, would postpone the development of ASI until proper mitigation efforts are developed and implemented (i.e., potentially decades). Practically speaking, this agreement is not going to be implemented tomorrow, but we hope it can serve as an end goal for AI governance, and in Appendix C we lay out a staged implementation that iteratively builds up to the final agreement.
Our proposed agreement is not without tradeoffs and risks: it forgoes beneficial AI capabilities, creates some pathways for authoritarian risk (although we think less risk than the current trajectory), cedes the U.S. technological lead in AI, risks leaking sensitive information due to verification, and would impose costs on industry and financial markets. Despite these downsides, we believe this agreement is worth pursuing because the alternative of racing towards ASI poses even greater risks.
@@
  • We provide a brief overview of AI risks, and an overview of the key details of the strategic situation informing our proposal.
@@
  • We provide the full text of an example agreement, detailed commentary, and connections to existing agreements in Appendix A. We discuss immediate steps different actors could take toward this agreement in Appendix B. We explain a staged implementation of the agreement that slowly builds from the current world to the agreement in Appendix C. We detail various approaches that could help locate existing AI chips and bring them under monitoring in Appendix D.
@@ 2 An overview of AI risks
There are many large-scale risks from advanced AI (i.e. AI much more powerful than today’s AI technology) that might warrant slowing or stopping AI progress for a significant period (e.g., decades). Most of these risks stem from future AIs becoming increasingly capable tools: Rapid automation may destabilize economies (drago2025intelligence; erdil2025gate). Capabilities that undermine strategic deterrence and incentivize first strikes may trigger wars and conflicts (PE-A3691-4; hendrycks_superintelligence_2025). Malicious actors may misuse AI to engage in wide-scale cyberattacks or to develop novel biological weapons (zelikow2024defense; crawford2024securing; esvelt2022delay). AI used for mass surveillance, propaganda, and police-state oppression may enable extreme concentrations of power (davidson2025aienabledcoupsho; bullock2025agi; drago2025intelligence).
@@
But while the above risks may lead to unprecedented catastrophes, humanity can probably survive them. A greater and more irreversible risk stems from AI systems that don’t remain passive tools, but instead become capable enough to permanently disempower humanity. This would plausibly lead to human extinction, either as an intentional act by the AIs (e.g., to prevent humans from making powerful competing AIs) or as a side effect of AIs optimizing for goals that don’t specifically designate human survival (yudkowsky2022agi; ngo_alignment_2025).
@@
This paper focuses on an international agreement to forestall ASI development until it can be done safely. A halt like this is primarily motivated by concerns about misalignment and AI takeover, and we think it is the only acceptable mitigation to those risks. But a halt would also be an effective tool for addressing many other risks, such as geopolitical destabilization, mass AI-caused unemployment, and misuse of powerful AIs (barnett2025ai). Therefore, others may support a halt even if they are not concerned with misalignment.
Our view is that AI developers are not on track to solve the alignment problem before the development of ASI. We will not fully defend this claim here—the difficulty of alignment is not this document’s central focus. Instead, we provide some general intuitions for why AI alignment appears challenging enough that success is far from guaranteed. This is sufficient to motivate our policy proposal.
@@
The deep learning paradigm underpinning modern AI development seems highly prone to producing agents that are not aligned with humanity’s interests. Under this paradigm, AI behavior is not hand-coded by human engineers, but instead emerges from automated training processes. Rather than producing lines of code that can be debugged when something goes wrong, these processes configure hundreds of billions to trillions of model weights that determine AI behavior. While it is sometimes possible to identify and activate features of these weights correlated with certain inputs or outputs, the deep machinery of these minds is effectively opaque to human observers (bereska2024mechanistic). No one can say with any certainty why an AI made a particular choice or how it would act in situations that differ from its training.
@@
Consequently, no one can manually adjust an AI’s weights to give it a particular goal or set of preferences. And yet, as seen in the growing ability of AIs to complete long-duration tasks, AIs are becoming increasingly capable of pursuing goals, e.g., in the setting of agentic software engineering (kwa2025measuring). Various well-documented examples of misbehavior by these same models (both in the lab and in the wild (baker2025monitoring; bondarenko2025demonstrating; lynch2025agentic; openai2025sycophancy; metr2025reward)) demonstrate that today’s training methods do not produce robustly aligned agents that could be trusted not to disempower humanity if they had the capacity.
@@
Many AI labs have a stated goal of building AIs that can surpass humans at all cognitive tasks. It is hard to predict when AI labs might succeed, but many predict ASI could be developed soon. The field has seen enormous progress in just five years—from AIs that couldn’t hold a conversation to “reasoning models” that solve hard exam questions (EpochLLMBenchmarkingHub2024) and that some major tech companies claim write a substantial fraction of their code (chandra2024ai; novet2025satya). Another few years of such progress might well push capabilities well beyond human level in all or nearly all domains.
@@
One accelerating factor complicating all forecasts is the possibility of AIs automating the process of AI research and development itself (davidson2025threetypesofinte; eth2025willairdautomati; wijk_re-bench_2024). The distance to this threshold is unclear, but some independent experts give median prediction dates ranging from 2028–2033 (lifland2025timelines). If or when this threshold is crossed, we may see a feedback loop where AI R&D is conducted by AIs that rapidly (in months or a few years) produce substantially superhuman AI (openai_openai_2025; anthropic_rsp_2025; google_frontier_safety_2025).
We note that artificial general intelligence (AGI)–systems that perform at human-level across all cognitive domains–would likely emerge before ASI. While AGI itself poses significant risks, our primary focus is preventing ASI, as misaligned superintelligence poses an irreversible existential threat. However, AGI capable of automating AI research could accelerate the path to ASI and trigger rapid, uncontrollable capability gains. This agreement aims to halt AI capabilities advancement before the creation of AGI systems that could automate AI research.
@@
On the current trajectory, the core challenge of AI alignment is unlikely to be adequately addressed before the first ASIs are built. Evidence for this can be seen in the failure to solve far easier related problems. Today’s frontier models are not robust to jailbreaks: It is relatively easy to craft inputs that lead to outputs that AI developers attempt to avoid (andriushchenko2024agentharm; andriushchenko2024jailbreaking; zou2023universal; anil_many-shot_2024; wang_jailbreak_2024). Current models, especially reasoning models, display a propensity to reward hack (exploiting unintended solutions) and lie to their users, against the best efforts of their developers (bondarenko2025demonstrating; wijk_re-bench_2024; chowdhury2025investigating; anthropic2025claude37; baker2025monitoring). Far tougher alignment challenges are likely to appear only at advanced capability levels where humanity would be unable to recover from failures (yudkowsky2022agi).
@@
There is a broad consensus in the field that catastrophic outcomes from misaligned AI pose a significant threat that should be addressed (bengio2024managing). Hundreds of luminaries signed on to a 2023 statement that “Mitigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war.” (cais_statement_2023) That same year, Yoshua Bengio, a Turing Award recipient and the most cited living scientist, said he thinks there is a “twenty percent probability that it turns out catastrophic” (lavoipierre2023ai). In a 2023 survey, 38% of survey respondents at top AI conferences said there was at least a 10% chance that AI’s outcome will be "extremely bad (e.g., human extinction)", conditional on the development of AI that can outperform humans on all tasks (grace2025thousands).
@@
This understanding extends to AI company leaders themselves: Anthropic CEO Dario Amodei has predicted a 10-25% that “something goes really quite catastrophically wrong on the scale of human civilization” (bartlett2023anthropic; amodei_axios_2025). OpenAI’s Sam Altman has called superintelligence “probably the greatest threat to the continued existence of humanity” (altman2015machine), and xAI’s Elon Musk has said, “AI is a fundamental risk to the existence of human civilization”  (london2017musk). Musk has indicated that his participation in this ASI race stems from the belief that it would happen even without him (musk2025interview). Those racing toward superintelligence, to the extent they are concerned with catastrophic risks, are stuck in a collective action problem (armstrong2016racing). Halting the race and avoiding extinction would require global coordination encompassing all relevant companies and governments.
@@ 3 The strategic situation
A 10% chance of extinction is unacceptable. To mitigate these risks, various approaches have been proposed: Have governments tightly regulate and license AI development (anderljung2023frontier; smith2024licensing), empower defensive actors early (buterin_my_2023; nielsen_notes_2024; bernardi_societal_2025), implement a joint global AI development project to avoid racing (brundage_my_2024; hausenloy_multinational_2023), race ahead while slowing opponents (aschenbrenner_situational_2024; amodei_machines_2024), or prevent the creation of ASI before we are ready (barnett2025ai; ramiah2025toward; aguirre2025keep).
@@
We think the only acceptable solution is to prevent the creation of ASI until catastrophic risks are mitigated. The state of AI alignment research is too nascent, and the problem too difficult. Nobody knows how to align advanced AIs to human values, including regulators or an international standards body (and we address some of these alternative plans in more depth in Section 5). Therefore, we think the right approach to this problem is to forestall the development of ASI for as long as is necessary to solve critical alignment challenges. We note that bans are not usually the best way to address dangerous technologies, but ASI is uniquely dangerous. An agreement establishing such a halt would need to be globally enforced, verifiable, and ready to implement once the political will to do so exists.
@@
For a halt to be effective, it must extend to all countries. If only cautious actors stop, other parties will continue the march toward superintelligence. The United States (U.S.) pausing on its own might buy around six months until the nearest company in the People’s Republic of China (PRC) catches up, and a U.S.–PRC agreement without global enforcement would similarly fall to the next closest actor, likely in a couple of years (artificialanalysis2025countries) (though it is unclear how long it would take to get from the current frontier to ASI). The U.S. and PRC are the preeminent superpowers and the centers of most frontier AI development; therefore, they are essential stakeholders in such a deal, and their strategic interests must be accommodated. Due to their influence, they are likely leaders in future AI agreements, whether or not other countries participate as well.
@@
A coalition including the U.S. and PRC would likely be able to attract nearly all other states, even if some are reluctant, given these two countries’ geopolitical influence and their current positions as AI leaders. We expect other countries to sign on to such an agreement once they understand AI risks better and conclude that international coordination is necessary to avoid these risks. Participation could also be encouraged with benefits such as security and access to AI infrastructure for allowed use. Where stronger incentives prove necessary, standard tools such as economic sanctions may apply. We believe a very ASI-concerned coalition headed by the U.S. and PRC would be sufficient to integrate other potential AI developers into the agreement or to sufficiently suppress other AI development so that it does not pose a threat.
@@
All parties to the agreement will desire assurance that other parties are not engaged in secret AI capabilities projects. Providing this assurance is a key aim of the agreement. Given a low-trust environment, parties must have the means to verify compliance to their satisfaction. It gives special consideration to the independent verification efforts of the U.S. and the PRC.
@@
There remain significant political obstacles to such an agreement occurring today. Nevertheless, to accept even a 10% chance of extinction is wholly inconsistent with how we manage other risks. We routinely adopt stringent mitigations for risks that are hundreds of times less likely and that would result in millions of times fewer fatalities, such as in the nuclear industry, and in many other fields (koessler2024risk; nrc2017approach; ehrhart2020evaluation).
@@ 4 The Agreement
AI development in the current deep learning paradigm is driven both by the large-scale employment of specialized computer chips and by improvements in AI algorithms (epoch2024trainingcomputeoffrontieraimodelsgrowsby45xperyear; kaplan_scaling_2020; ho_algorithmic_2024). Therefore, the coalition stops the advancement of AI capabilities toward superintelligence by restricting the scale of AI training (operationalized as total training FLOP) and by restricting research into AI algorithms and other areas that would undermine the governance regime. These research restrictions are narrowly targeted at research that advances towards ASI or would undermine the agreement.
AI training runs above the Strict Threshold (i.e., $10^{24}$ FLOP) are prohibited. Training runs below this threshold but above the Monitored Threshold (i.e., $10^{22}$ FLOP) must be approved and monitored by coalition authorities. Training runs below the Monitored Threshold require no approval or monitoring.
There are numerous considerations in deciding where these thresholds should be set, and these thresholds can be modified during agreement negotiations or after the agreement is in effect. First, nobody knows what scale of training would reach particularly dangerous AI capability levels; given this uncertainty and the massive negative effect of setting thresholds too high, we suggest a conservative approach. Second, many near-frontier AI models of today are trained at a scale only slightly above the Strict Threshold; for example, DeepSeek-R1 ($\sim 4\times 10^{24}$ FLOP) and gpt-oss-120B ($\sim 5\times 10^{24}$ FLOP) (EpochAIModels2025). See Figure 2 for a visualization of how these thresholds relate to the training scale of existing AI models. Due to improvements in AI algorithms and data, the capability of models trained at a given computational scale increases rapidly over time (ho_algorithmic_2024). Due to likely progress in algorithms and data between today and when this agreement would come into effect, AIs trained at the Strict Threshold will be more capable—potentially much more—than the models trained at that scale today. There are numerous other considerations in what these thresholds should be and whether FLOP thresholds are appropriate (erben_JRC143255; hooker2024limitations; heim_training_2024); a full discussion is out of scope for this paper.
@@
Verification of these restrictions is practical because AI chips are expensive and specialized, and thousands of them are needed for frontier AI development. This approach helps ensure that training runs that advance toward ASI are prevented while still allowing continued use of current AI technology.
@@
Figure 2: The training compute used to train various notable AI models in the last few years, along with the two thresholds in this agreement. In the agreement, new training above the Strict Threshold ($10^{24}$ FLOP) would be prohibited and new training between the Monitored Threshold ($10^{22}$ FLOP) and Strict Threshold ($10^{24}$ FLOP) would be monitored. Due to a lack of public data, only some models have confident estimates for the training compute used. The number in parentheses is the model’s score on the Artificial Analysis Intelligence Index. Data is based on EpochAIModels2025 and artificialanalysis_models. Cost estimates are based on October 2025 rental prices for B200 GPUs from CoreWeave, assuming 50% utilization in FP8.
@@
To achieve this verification, the coalition locates existing AI chips, tracks new AI chips being produced, and monitors AI chip use in order to ensure chips are not being used for restricted activities (scher_mechanisms_2024; baker2025verifying; harack2025verification). Chips are located through coordinated efforts by intelligence agencies and domestic law-enforcement agencies from coalition countries. These efforts involve combining supply chain tracking, mandatory reporting, state intelligence gathering, open-source intelligence, power consumption monitoring, challenge inspections, whistleblowers, and more. Appendix D provides more detail on how this could be done. Tracking new AI chips relies on the concentrated chip supply chain and serves the goal of making sure chips are not smuggled outside of the monitoring regime. Finally, the chips themselves are monitored in order to verify that they are only being used for acceptable use cases.
@@
AI chip monitoring starts with discerning whether the chips are being used for inference on existing AIs or training of new AIs. The coalition works to develop tamper-resistant on-chip mechanisms for such purposes, as finer-grained control permits greater use of AI chips for safe applications, with fewer inspections. Initially inspectors are given ongoing physical access to chips, as is likely needed for robust chip-use verification (aarne2024secure; kulp2024hardwareenabled). To ensure chip use verification is applied, the coalition prohibits large concentrations of chips (i.e., greater than 16 H100-equivalents; 16 H100s cost approximately \$500,000 USD in 2025) outside of monitored facilities.
@@
Despite prohibitions on large-scale training, AIs could continue getting more capable via improvements to the algorithms and software used to train them. Therefore, the coalition adopts appropriate restrictions on research that contributes to frontier AI development or research that could endanger the verification methods in the agreement. These restrictions would cover certain machine learning-related research and may eventually expand to include other AI paradigms if those paradigms seem likely to lead to ASI. Coalition members draw on their experience restricting research in other dangerous fields such as nuclear and chemical weapons. These research restrictions aim to be as narrow as possible, preventing the creation of more capable general AI models while still allowing safe and beneficial narrow AI models to be created. The coalition encourages and makes explicit carve-outs for safe, application-specific AI activities, such as self-driving cars and other uses that provide benefits to society.
@@
Verification that members are adhering to research restrictions is aided by the fact that relatively few people have the skills to contribute to this research—likely only thousands or tens of thousands, see Appendix A, Article IX. Nations verify compliance by way of intelligence gathering, interviews with the researchers, whistleblower programs, and more. This verification aims to use non-invasive measures to ensure that researchers are not working on restricted topics. Additionally, inspectors verify that the medium-scale training allowed by the agreement uses only approved methods (e.g., from a Whitelist) and does not make use of any novel AI methods or algorithms; these would be evidence that restricted research had taken place.
@@
Should nonproliferation fail, or a rogue actor is found to be working towards ASI, coalition members can employ their agreed upon enforcement mechanisms. These start with standard tools of international pressure such as diplomacy and economic sanctions. Depending on the nature of the developments, countries could escalate to other means at their disposal to disrupt, slow, or impede ASI development. This draws parallels to other contexts where countries have intervened to stop violations of arms-control and nonproliferation agreements.
@@
While the goal is for this agreement to last as long as is necessary, that may not be practical. There are numerous reasons the agreement might fail; for instance, two threat actors are countries outside the agreement and covert state-backed ASI efforts from countries that are in the agreement. The verification, nonproliferation, and enforcement measures in the agreement are set up with the purpose of preventing such failures, but it’s hard to predict how successful they will be. World leaders may find themselves weighing the risk of a covert project building ASI despite the agreement against the risk of nullifying the agreement and building ASI themselves.